show clear Part Il Short Answer Questions (36 pts.) (Write legible and coherent anawer s sequence ofkey calculations or considerations that led to the final anowers) S1) (3 pts.) Given the following equilibria, NF(aq) + 2 OH (aq) Ni(aq) + 4 CN (aq) Ni(OH)2(s) Ni(CN)-(aq) K-18 x 10 K3-2.0x 10 Plus determine the equilibrium constant, Ke for the following reaction Ni(OH) (s) +4 CN (aq)Ni(CN) (aq)+ 2 OH (aq) S2) Consider the following equilibrium: Equal amounts of 0.180 moles of H:(g) and CO(g) are placed in a 1.0 L container and heated to 1260 a 3 pts.) If Ke 1.6 at 1260K, what is the equilibrium concentration of CO(g? b) (2 pts.) At another temperature, it was found that the amount of Ha(g) has decreased to 9.036 mo reaction system reaches equilibrium. What is the equilibrium constant, Ks, for this tcmperature?
Answer #1

S1) FROM the given data,

     revesre eq1 + eq2 = desired equation.

     so that,

     (1/k1)*(k2) = Ke

Ke = (1/(1.8*10^25))*(2.0*10^28)

     = 1.11*10^3

S2)

a)        CO2(g)   + H2(g) <=====> CO(g) + H2O(g)

initial 0.18 M     0.18 M           -       -

change    -x         -x              x        x

equil   0.18-x     0.18-x            x        x

   Kc = [CO][H2O]/[CO2][H2]

1260 = (X^2/((0.18-X)^2)

   X = 0.175

AT equilibrium,

[CO2] = 0.18-0.175 = 0.005 M

[H2] = 0.18-0.175 = 0.005 M

[CO] = X = 0.175 M

[H2O] = X = 0.175 M

b)       CO2(g)   + H2(g) <=====> CO(g) + H2O(g)

initial 0.18 M     0.18 M           -       -

change   -0.144 M   -0.144 M       +0.144 M +0.144 M

equil   0.036     0.036           0.144 M   0.144 M

   Kc = [CO][H2O]/[CO2][H2]

       = (0.144*0.144)/(0.036*0.036)

       = 16